OpenStudy (rubylovesjaune):

Solve the formula for the variable y. z - 4x + 2y

OpenStudy (ashy98):

z=4x+2y -4x from each side z-4x = 2y divide each side by 2 \[\frac{ z-4x }{ 2 }\]

OpenStudy (rubylovesjaune):

Okay thank you.

OpenStudy (ashy98):

Np

OpenStudy (erak):

Well it's: 0 = z - 4x + 2y 2y = 4x - z y = 2x - z/2 @Ashy98 Your "2y" should be negative

OpenStudy (mathmale):

I have a problem with this discussion because you don't have an equation here. You cannot solve an expression for a given letter when it's not an equation. Have you copied down the problem statement correctly?

OpenStudy (ashy98):

My bad thank you for seeing that !@Erak